OpenStudy (anonymous):

find the domain of the function square root of 6x-42

sam (.sam.):

Find the restrictions of x, ie when \(\sqrt{-x}\)

sam (.sam.):

\[\sqrt{6x-42}\] Or just solve 6x-42=0 x=7 So if x is larger than 7, then we'll have all positive values

OpenStudy (anonymous):

i understand that part i just dont know how to write the answer in correct form

sam (.sam.):

x equal to 7 will work as well, it'll give \(\sqrt{0}\)

sam (.sam.):

Write \[\left\{ x \in \mathbb{R} : x \geq 7 \right\}\]

OpenStudy (anonymous):

do i use brackets or parenthesis? and is it just (x> or = 7)

OpenStudy (anonymous):

the true values of the range are considered to be b/w 0 and infinit (for which i wlii write only q) so let the given func be y 0<y<q 0< (6x-42)^(1/2)< q suqaring the iniquality 0< (6x- 42)< q adding 42 and dividing by 6 the whole iniquility 42/6 < x < q thus 7< x< q

OpenStudy (anonymous):

so \[x \in [7,\infty)\]

sam (.sam.):

Use { }